Can you show me the steps for solving a system of equations
3x-4=7
Y+2x=8
Thank you

Answer by Alan3354(69443)

Can you show me the steps for solving a system of equations
3x-4=7 --> 3x = 11 --> x = 11/3
Y+2x=8
-----
y + 2x = 8
y + 2*(11/3) = 8
y = 8 - 22/3
y = 2/3